# references.bib

### v1

@@ 2.2 Application of AI in Art Creation
With the rapid development of AI technology, the application of deep learning models in art creation has attracted increasing attention, particularly the convolutional neural network (CNN)-driven neural style migration technique ([ref8, ref9, ref10]). The neural style migration technique enables the generation of images with different artistic styles by extracting the features of content images and style images. The VGG19 model, a pre-trained deep convolutional network, is a popular choice for advanced feature extraction. In 2016, Leon Gatys et al. proposed a neural style migration method based on the VGG19, which generates a high-quality style through the optimisation of the content loss and the style loss migration images ([ref11, ref12]). This method outperforms Generative Adversarial Networks (GANs) in controlling the generated details and maintaining content consistency, and thus becomes an ideal tool for artistic image processing and is widely used in art creation and digital style simulation.
@@ 2.3 Physical Properties in Works of Art
In 2024, Yinxiang et al. demonstrated that the vortex flow characteristics observed in Van Gogh’s Night of the Stars and Moon are highly compatible with the phenomenon of turbulence in physics, particularly in accordance with Kolmogorov’s theory of turbulence ([ref13]). In accordance with the theory, energy is transferred from large-scale vortices to small-scale vortices in a gradual manner, and ultimately dissipated. It was observed that the vortices present in the work not only align with the Kolmogorov turbulence scalar law, but also with the scalar turbulence power spectrum proposed by George Batchelor in 1959.
@@ 3.4.1 Feature Hashing
In order to guarantee the uniqueness and integrity of the generated digital artefacts, this study employs a methodology that combines multiple fractal spectra with turbulence features to create unique identifiers, designated as Token IDs. In particular, the fractal dimensions and turbulence features are initially integrated into a unified feature vector, which serves as the fundamental descriptive information pertaining to the artwork. Subsequently, the aforementioned feature vector is hashed using cryptographic hashing algorithms, such as SHA-256 ([ref24]), in order to generate a unique Token ID. This not only encodes the distinctive features of the artwork in question, but also provides a means of verifying the artwork’s uniqueness on the blockchain, thereby preventing any instances of forgery or tampering. This mechanism serves to guarantee the authenticity and transparency of digital artworks, whilst simultaneously supporting their traceability and trading in a decentralised network. This, in turn, serves to enhance the trust and value of artworks in the NFT market.
@@ 3.4.2 NFT Metadata Generation
The generated Token IDs and their corresponding feature data are integrated into the NFT metadata, thereby ensuring authentication and traceability of digital artworks. The metadata is structured as follows:

### v3

@@ 2.2 Application of AI in Art Creation
With the rapid development of AI technology, the application of deep learning models in art creation has attracted increasing attention, particularly the convolutional neural network (CNN)-driven neural style migration technique ([ref8, ref9, ref10]). The neural style migration technique enables the generation of images with different artistic styles by extracting the features of content images and style images. The VGG19 model, a pre-trained deep convolutional network, is a popular choice for advanced feature extraction. In 2016, Leon Gatys et al. proposed a neural style migration method based on the VGG19, which generates a high-quality style through the optimisation of the content loss and the style loss migration images ([ref11, ref13]). This method outperforms Generative Adversarial Networks (GANs) in controlling the generated details and maintaining content consistency, and thus becomes an ideal tool for artistic image processing and is widely used in art creation and digital style simulation.
@@ 2.3 Physical Properties in Works of Art
In 2024, Yinxiang et al. demonstrated that the vortex flow characteristics observed in Van Gogh’s Night of the Stars and Moon are highly compatible with the phenomenon of turbulence in physics, particularly in accordance with Kolmogorov’s theory of turbulence ([ref12]). In accordance with the theory, energy is transferred from large-scale vortices to small-scale vortices in a gradual manner, and ultimately dissipated. It was observed that the vortices present in the work not only align with the Kolmogorov turbulence scalar law, but also with the scalar turbulence power spectrum proposed by George Batchelor in 1959.
@@ 3.4.1 Feature Hashing
In order to ensure the uniqueness and integrity of the generated digital artifacts, this study employs a methodology that combines multiple fractal spectra with turbulence features to create unique identifiers, referred to as Token IDs. The process is as follows:First, the fractal dimensions and turbulence features are integrated into a unified feature vector. The fractal dimension quantifies the self-similarity of an image, reflecting its fractal characteristics and highlighting the details and layers of the artwork. Turbulence features, on the other hand, are statistical metrics that describe the dynamic complexity of an image, capturing its fluidity and intricate structure. Together, these features form the fundamental descriptive information pertaining to the artwork.Next, the aforementioned feature vector is hashed using cryptographic hashing algorithms, such as (SHA-256 [ref24]), to generate a unique Token ID. The hashing algorithm transforms the input feature vector into a fixed-length hash value, ensuring the uniqueness and irreversibility of each Token ID.
@@ 3.4.2 NFT Metadata Generation
The generated Token IDs and their corresponding feature data are integrated into the NFT metadata, ensuring the authentication and traceability of digital artworks. The metadata is structured as follows: